OpenStudy (anonymous):

Find the distance d(P1, P2) between the points P1 and P2. P1=(-5,1) P2=3,3) d(P1, P2)=

OpenStudy (anonymous):

you can plot the points to make a line then the path from (-5,1) and (3,3) would for a triangle so you can use pytagorean theorem for it so it would be 8^2 + 2^2 = c^2 64 + 4 = c^2 c^2 = 68 c = sqrt(68) <--- distance

OpenStudy (anonymous):

Ok, here is as far as I can get....d=\[\sqrt{(x _{2}-x _{1})^2+(y _{2}-y _{1})^{2}} \] \[x ^{1}=-5, y ^{1}=1, x ^{2}=3, y ^{2}=3\]

OpenStudy (anonymous):

you can just plug in the points then x2 - x1 = 8 and y2 - y = 2

OpenStudy (anonymous):

I still do not get it

OpenStudy (anonymous):

I am not the brightest in this subject...sorry :(